1. A basketball player can jump 0.72m off the floor from a standing position. Calculate the player’s speed when leaving the floor
1
Expert's answer
2020-12-06T17:16:29-0500

The potential energy of the basketball player at the highest point is:


"U = mgh"

where "m" is the mass of the player, "h = 0.72m" is the height and "g = 9.81m\/s^2" is the gravitational acceleration. Using the energy conservation law find, that this potential energy is equal to the kinetic energy when player was leaving the floor:


"U = K=\dfrac{mv^2}{2}"

where "v" is the player’s speed when leaving the floor. Substituting the expression for "U" and expressing "v", obtain:


"mgh=\dfrac{mv^2}{2}\\
v = \sqrt{2gh}\\
v = \sqrt{2\cdot 9.81\cdot 0.72} \approx 3.76m/s"

Answer. 3.76 m/s.